A probability experiment consists of rolling a fair 12​-sided die. Find the probability of the event below. rolling a 5
Answer #1

There are 12 possible outcomes when rolling 12-sided die once and these outcomes are as follows:

S = { 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12 }

Therefore, probability of rolling a 5 = 1/12 = 0.0833

Probability = 1/12 = 0.0833